# Batch Calibration: Rethinking Calibration for In-Context Learning and Prompt Engineering

### v2

@@ 2 Related Work
Lu et al. (2022) show the sensitivity of LLMs to ICL examples. This phenomenon is further explained through the effect from pretraining term frequencies (Razeghi et al., 2022) and corpora (Shin et al., 2022). Meanwhile, Xie et al. (2022) explain the ICL process through implicit Bayesian inference, and Wei et al. (2023) show the emergent ability of LLMs by learning new input-label mappings. Various methods have been proposed to optimally select better in-context templates (Sorensen et al., 2022; Pan et al., 2023; Yin et al., 2023) and examples (Rubin et al., 2022; Liu et al., 2022a; Wan et al., 2023b). Specifically, Wan et al. (2023a) introduce a selection criteria based on the consistency, diversity, and repetition of in-context examples. Recently, noisy channel prompting (Min et al., 2022a) and flipped learning (Ye et al., 2023) have been proposed for robust ICL. Learning to assign labels by k-nearest neighbors (Xu et al., 2023) and training decoder networks (Cui et al., 2023) are also effective alternatives for few-shot ICL.
@@ 3.3 Design Principles Behind Calibrations
Method Token #Forward Comp. Cost Cali. Form Learning Term Decision Boundary $h(\mathbf{p})$ Multi- Sentence Multi- Class
CC N/A
$1+1$
Inverse $\mathbf{W}\mathbf{p}+\mathbf{b}$ $\mathbf{W}=\operatorname*{diag}(\hat{\mathbf{p}})^{-1}$, $\mathbf{b}=\mathbf{0}$ $p_{0}=\alpha p_{1}$
DC Random
$20+1$
Add $\mathbf{W}\mathbf{p}+\mathbf{b}$ $\mathbf{W}=\mathbf{I}$, $\mathbf{b}=-\frac{1}{T}\sum_{t}\mathbf{p}(y|\text{text}_{j},C)$ $p_{0}=p_{1}+\alpha$
PC -
$1$
EM-GMM - $\sum_{j}\alpha_{j}P_{G}(\mathbf{p}|\mathbf{\boldsymbol{\mu}_{j}},\mathbf{\boldsymbol{\Sigma}_{j}})$ $P_{\text{G}}(\mathbf{p}|\mu_{0},\Sigma_{0})=P_{\text{G}}(\mathbf{p}|\mu_{1},\Sigma_{1})$
BC (Ours) -
$1$
Add $\mathbf{W}\mathbf{p}+\mathbf{b}$ $\mathbf{W}=\mathbf{I}$, $\mathbf{b}=-\operatorname*{\mathbb{E}}_{x}\Big{[}\mathbf{p}(y|x,C)\Big{]}$ $p_{0}=p_{1}+\alpha$
@@
To address this research question, we first derive the decision boundary for each category of calibration methods. We recall that the classification by a LLM is based on $\operatorname*{arg\,max}_{j\in\{0,...,J-1\}}p_{j}$ where $p_{j}$ denotes the $j$-th element of output vector $\mathbf{p}$. Consider binary classification problem for simplicity: the decision boundary $h(\mathbf{p})$ for ICL is given by the line $p_{0}-p_{1}=0$: the model predicts class 0, $y_{0}$, if $p_{0}-p_{1}\geq 0$, and class 1 otherwise. Consequently, CC and DC that apply an affine transformation at $\mathbf{p}$ is equivalent to a linear transformation to the decision boundary. In CC with $\mathbf{W}=\operatorname*{diag}(\hat{\mathbf{p}})^{-1}$, $\mathbf{b}=\mathbf{0}$, the decision boundary can then be derived as $p_{0}\times\frac{1}{\hat{p}_{0}}=p_{1}\times\frac{1}{\hat{p}_{1}}\rightarrow p_{0}-p_{1}\times\frac{\hat{p}_{0}}{\hat{p}_{1}}=0$, which is a rotation of the ICL’s linear decision boundary around the origin. Similarly, DC with $\mathbf{W}=\mathbf{I}$, $\mathbf{b}=-\frac{1}{T}\sum_{t}\mathbf{p}(y|[\textsc{Random Text}]_{t},C)=-\hat{\mathbf{p}}$ is equivalent to a shift of ICL’s linear decision boundary away from the origin, such that $p_{0}-p_{1}=(\hat{p}_{0}-\hat{p}_{1})$. It is worth noting that both calibration choices lead to a linear decision boundary, indicating that the calibration problem can be framed as an unsupervised decision boundary learning problem. Based on this intuition, we further derive the decision boundary for PC as $P_{\text{G}}(\mathbf{p}|\mu_{0},\Sigma_{0})-P_{\text{G}}(\mathbf{p}|\mu_{1},\Sigma_{1})=0$, which delivers a non-linear boundary between the estimated Gaussian mixtures. We conduct a preliminary experiment to visualize the derived decision bounds from existing calibration methods alongside the ICL baseline. In Fig. 2, we observe that uncalibrated ICL is biased towards predicting negative in the SST-2 task. This biased prediction is then mitigated by each calibration method, where we observe a rotated decision boundary from CC, a shifted boundary from DC, and a non-linear boundary between the GMMs by PC. However, in the QNLI task (bottom row of Fig. 2), we observe failure cases in the calibration baselines, in particular, PC (third figure from the left), where it fails to capture the correct distribution for each class. From Fig. 2 and the additional results in Fig. 10 in Appendix §E, we find that while theoretically more flexible, the non-linear decision boundaries learned by PC tend to be susceptible to overfitting and may suffer from instability in EM-GMM. We hypothesize that the PC boundary is even more vulnerable to instability for more challenging multi-class tasks due to the increased difficulties of learning clusters and assigning classes correctly. Conversely, we find that linear decision boundaries, as evidenced by CC and DC, can be more robust and generalizable across tasks. We validate this hypothesis by proposing BC with extensive experiments in Sec. 5.2.
@@ 4 Batch Calibration
Model
PaLM 2-S PaLM 2-L
Method
ICL CC DC PC BC ICL CC DC PC BC
SST-2
93.620.62 95.500.25 94.290.32 95.710.10 95.440.15 93.165.18 95.820.62 94.912.01 95.640.47 95.780.55
MNLI
68.527.98 60.0711.26 63.451.99 59.2913.79 75.122.76 72.773.65 79.453.46 71.534.86 78.687.10 81.342.29
QNLI
81.201.90 56.863.29 65.623.53 69.8217.73 82.451.82 64.683.53 69.714.89 68.973.27 61.0115.26 87.901.24
MRPC
66.4210.15 70.440.94 68.580.21 71.861.29 70.052.40 73.191.21 72.403.53 68.680.40 75.392.60 70.392.56
QQP
63.910.66 65.555.34 53.929.35 65.283.42 71.481.46 82.570.75 81.172.03 78.321.82 81.420.24 79.561.40
BoolQ
83.993.90 87.141.60 87.641.10 88.700.15 87.830.10 90.020.60 90.150.54 87.771.17 64.4022.37 90.100.22
CB
45.7110.61 29.647.85 65.713.20 81.079.42 78.213.19 92.862.19 85.727.78 92.862.82 89.297.25 93.211.49
COPA
96.402.30 95.802.05 96.402.88 96.202.05 96.402.07 98.601.14 97.201.10 97.400.89 99.000.71 97.001.00
RTE
80.941.29 79.780.92 76.821.72 80.431.07 83.471.10 75.092.11 80.002.48 79.211.95 86.642.62 85.422.48
WiC
50.690.59 50.560.50 49.970.13 51.383.56 61.102.07 51.351.90 55.586.38 54.676.02 57.8711.08 64.838.59
ANLI-R1
46.244.21 42.543.20 40.263.66 40.286.46 59.820.51 63.062.63 71.923.71 73.563.88 72.308.05 75.003.03
ANLI-R2
40.440.90 38.360.82 38.443.46 41.884.50 50.160.82 58.401.19 65.363.75 65.481.91 64.982.94 67.302.34
ANLI-R3
42.530.99 38.781.04 43.675.25 37.500.81 55.751.66 61.353.14 67.320.98 66.230.72 63.036.03 66.380.74
Avg.
66.20 62.39 64.98 67.65 74.41 75.16 77.83 76.89 76.13 81.09
@@ 5.2 Main Experiments
We further analyze the performance of BC by varying the ICL shots from 0 to 4 shots as shown in Fig. 6, and BC again outperforms all baseline methods. We also observe an overall trend for improved performance when more shots are available, and the performance disparities between BC and ICL converge on some tasks, which suggests that BC allows LLMs to more effectively take advantage of more in-context demonstrations.
We also observe that PC exhibits the worst stability across the number of shots. In Table 3, we show that further slight gains, 1% on average, can be achieved by involving an adjustable strength parameter that refines the calibration and mitigates estimation errors. This alternative design not only makes BC applicable to both zero-shot and few-shot setups but also shows its capability to improve further from limited labeled data.
Method
SST-2 MNLI
QNLI
MRPC QQP BoolQ CB
COPA
RTE WiC ANLIR1 ANLIR2 ANLIR3 Avg.
BC
95.4 75.0
83.5
68.6 70.3 87.9 75.0
98.0
84.1 63.3 59.8 51.1 53.3 74.3
BCL
96.3 75.0 83.5 74.3 72.3 88.8 83.9 99.0 82.7 63.2 58.0 49.7 52.2 75.3
@@ 5.3 Robustness and Ablation Studies
We analyze the robustness of BC with respect to common prompt engineering design choices that were previously shown to significantly affect LLM performance (Lu et al., 2022; Liu et al., 2022a): choices and orders of in-context examples, the prompt template for ICL, and the verbalizers, as shown in Fig. 9 evaluated on RTE. Setup details are listed in Appendix §F. First, we find that BC is more robust to ICL choices and can mostly achieve the same performance with different ICL examples. Additionally, given a single set of ICL shots, altering the order between each ICL example has minimal impact on the BC performance. However, it is worth noting that an optimal order selection can still lead to promising ICL performance. Furthermore, we analyze the robustness of BC under 10 designs of prompt templates, where BC shows consistent improvement over the ICL baseline. Therefore, though BC makes further improvements, a well-designed template can further enhance the performance of BC. Lastly, we examine the robustness of BC to variations in verbalizer designs. Remarkably, even when employing unconventional choices such as emoji pairs as the verbalizers (listed in Tables 7 & F) leading to dramatic oscillations of ICL performance, BC largely recovers performance. This observation shows BC robustifies LLM predictions under common prompt design choices and makes prompt engineering easier.
@@ Appendix A Additional Related Work
Prompt learning is an efficient learning pipeline for LLM as an alternative to traditional full-model fine-tuning (Liu et al., 2023). Soft prompting (Li & Liang, 2021; Lester et al., 2021; Liu et al., 2022b) enables fast adaptation of LLM by appending learnable continuous prompts in the embedding space while freezing the rest of weights. The recent development of parameter-efficient fine-tuning methods (Houlsby et al., 2019; Hu et al., 2022; Zhou et al., 2023a), which learn additional modules, may also be interpreted as a form of soft prompting (He et al., 2022). However, these soft prompt learning methods inevitably require gradients and internal model access. On the other hand, hard prompting (Shin et al., 2020) is an appealing learning category for learning discrete prompts. Recent efforts have been devoted to black-box prompt search without accessing model gradients, and more interpretable prompts can be found by reinforcement learning (Deng et al., 2022; Zhang et al., 2023), gradient estimation (Diao et al., 2023), and other derivative-free search algorithms (Prasad et al., 2023; Zhou et al., 2023b).
@@
Test-time adaptation aims to mitigate the domain covariate shift using the test-time statistics. Wang et al. propose TENT that minimizes the entropy by updating the affine parameters in the BN layer. Nado et al. (2020) and Schneider et al. (2020) introduce using test-time batch statistics for the standardization in BN and mixing it with source statistics to conquer covariate shift, respectively. Similarly, mixing the statistics with predefined hyperparameters (You et al., 2021; Khurana et al., 2021), interpolating source and target-domain statistics (Lim et al., 2023), or using a running average estimate (Mirza et al., 2022; Hu et al., 2021) have also been proposed to adapt the BN layer. Zou et al. (2022) introduce strength parameters in adapting the standardization statistics in semantic segmentation tasks. We differentiate from test-time BN approaches by mitigating the bias in the novel context of LLM, and there is no source statistic similar to a BN layer in computer vision backbones.
@@ Appendix C Dataset Statistics
Dataset Objective #sentences #classes |Test|

### v3

@@ 2 Related Work
Lu et al. (2022) show the sensitivity of LLMs to ICL examples. This phenomenon is further explained through the effect from pretraining term frequencies (Razeghi et al., 2022) and corpora (Shin et al., 2022). Meanwhile, Xie et al. (2022) explain the ICL process through implicit Bayesian inference, and Wei et al. (2023) show the emergent ability of LLMs by learning new input-label mappings. Various methods have been proposed to optimally select better in-context templates (Sorensen et al., 2022; Pan et al., 2023; Yin et al., 2023) and examples (Rubin et al., 2022; Liu et al., 2022a; Wan et al., 2023b). Specifically, Wan et al. (2023a) introduce a selection criteria for in-context examples based on the confidence. Recently, noisy channel prompting (Min et al., 2022a), flipped learning (Ye et al., 2023), and self-ensembling (Li et al., 2023) have been proposed for robust ICL. Learning to assign labels by k-nearest neighbors (Xu et al., 2023) and training decoder networks (Cui et al., 2023) are also effective alternatives for few-shot ICL.
@@ 3.1 Bias in Prompting and In-Context Learning (ICL)
{tblr}
column2-9 = c, vline2 = -, hline1-2,6 = -, hline1,6 = -0.08em, stretch = 0.5 Method & Token #Forward Comp. Cost Cali. Form Learning Term Decision Boundary $h(\mathbf{p})$ Multi- Sentence Multi- Class
CC N/A $1+1$ Inverse $\mathbf{W}\mathbf{p}+\mathbf{b}$ $\mathbf{W}=\operatorname*{diag}(\hat{\mathbf{p}})^{-1}$, $\mathbf{b}=\mathbf{0}$ $p_{0}=\alpha p_{1}$
DC Random $20+1$ Add $\mathbf{W}\mathbf{p}+\mathbf{b}$ $\mathbf{W}=\mathbf{I}$, $\mathbf{b}=-\frac{1}{T}\sum_{t}\mathbf{p}(y|\text{text}_{j},C)$ $p_{0}=p_{1}+\alpha$
PC - $1$ EM-GMM - $\sum_{j}\alpha_{j}P_{G}(\mathbf{p}|\mathbf{\boldsymbol{\mu}_{j}},\mathbf{\boldsymbol{\Sigma}_{j}})$ $P_{\text{G}}(\mathbf{p}|\mu_{0},\Sigma_{0})=P_{\text{G}}(\mathbf{p}|\mu_{1},\Sigma_{1})$
BC (Ours) - $1$ Add $\mathbf{W}\mathbf{p}+\mathbf{b}$ $\mathbf{W}=\mathbf{I}$, $\mathbf{b}=-\operatorname*{\mathbb{E}}_{x}\Big{[}\mathbf{p}(y|x,C)\Big{]}$ $p_{0}=p_{1}+\alpha$
@@ 3.3 Design Principles Behind Calibrations
To address this research question, we first derive the decision boundary for each category of calibration methods. We recall that the classification by a LLM is based on $\operatorname*{arg\,max}_{j\in\{0,...,J-1\}}p_{j}$ where $p_{j}$ denotes the $j$-th element of output vector $\mathbf{p}$. Consider binary classification problem for simplicity: the decision boundary $h(\mathbf{p})$ for ICL is given by the line $p_{0}-p_{1}=0$: the model predicts class 0, $y_{0}$, if $p_{0}-p_{1}\geq 0$, and class 1 otherwise. Consequently, CC and DC that apply an affine transformation at $\mathbf{p}$ is equivalent to a linear transformation to the decision boundary. In CC with $\mathbf{W}=\operatorname*{diag}(\hat{\mathbf{p}})^{-1}$, $\mathbf{b}=\mathbf{0}$, the decision boundary can then be derived as $p_{0}\times\frac{1}{\hat{p}_{0}}=p_{1}\times\frac{1}{\hat{p}_{1}}\rightarrow p_{0}-p_{1}\times\frac{\hat{p}_{0}}{\hat{p}_{1}}=0$, which is a rotation of the ICL’s linear decision boundary around the origin. Similarly, DC with $\mathbf{W}=\mathbf{I}$, $\mathbf{b}=-\frac{1}{T}\sum_{t}\mathbf{p}(y|[\textsc{Random Text}]_{t},C)=-\hat{\mathbf{p}}$ is equivalent to a shift of ICL’s linear decision boundary away from the origin, such that $p_{0}-p_{1}=(\hat{p}_{0}-\hat{p}_{1})$. It is worth noting that both calibration choices lead to a linear decision boundary, indicating that the calibration problem can be framed as an unsupervised decision boundary learning problem. Based on this intuition, we further derive the decision boundary for PC as $P_{\text{G}}(\mathbf{p}|\mu_{0},\Sigma_{0})-P_{\text{G}}(\mathbf{p}|\mu_{1},\Sigma_{1})=0$, which delivers a non-linear boundary between the estimated Gaussian mixtures. We conduct a preliminary experiment to visualize the derived decision bounds from existing calibration methods alongside the ICL baseline. In Fig. 2, we observe that uncalibrated ICL is biased towards predicting negative in the SST-2 task. This biased prediction is then mitigated by each calibration method, where we observe a rotated decision boundary from CC, a shifted boundary from DC, and a non-linear boundary between the GMMs by PC. However, in the QNLI task (bottom row of Fig. 2), we observe failure cases in the calibration baselines, in particular, PC (third figure from the left), where it fails to capture the correct distribution for each class. From Fig. 2 and the additional results in Fig. 10 in Appendix §E, we find that while theoretically more flexible, the non-linear decision boundaries learned by PC tend to be susceptible to overfitting and may suffer from instability in EM-GMM.
We hypothesize that the PC boundary is even more vulnerable to instability for more challenging multi-class tasks due to the increased difficulties of learning clusters and assigning classes correctly. Conversely, we find that linear decision boundaries, as evidenced by CC and DC, can be more robust and generalizable across tasks. We validate this hypothesis by proposing BC with extensive experiments in Sec. 3.
@@ 4 Batch Calibration
{tblr}
width = cell12 = c=5c, cell17 = c=5c, cell22-11 = c=1, vline2-3 = 1, vline2,7 = 1-17, hline1-3,16-17 = -, hline1,17 = -0.08em, stretch = 0.5 Model & PaLM 2-S PaLM 2-L
Method ICL CC DC PC BC ICL CC DC PC BC
SST-2 93.620.62 95.500.25 94.290.32 95.710.10 95.440.15 93.165.18 95.820.62 94.912.01 95.640.47 95.780.55
MNLI 68.527.98 60.0711.26 63.451.99 59.2913.79 75.122.76 72.773.65 79.453.46 71.534.86 78.687.10 81.342.29
QNLI 81.201.90 56.863.29 65.623.53 69.8217.73 82.451.82 64.683.53 69.714.89 68.973.27 61.0115.26 87.901.24
MRPC 66.4210.15 70.440.94 68.580.21 71.861.29 70.052.40 73.191.21 72.403.53 68.680.40 75.392.60 70.392.56
QQP 63.910.66 65.555.34 53.929.35 65.283.42 71.481.46 82.570.75 81.172.03 78.321.82 81.420.24 79.561.40
BoolQ 83.993.90 87.141.60 87.641.10 88.700.15 87.830.10 90.020.60 90.150.54 87.771.17 64.4022.37 90.100.22
CB 45.7110.61 29.647.85 65.713.20 81.079.42 78.213.19 92.862.19 85.727.78 92.862.82 89.297.25 93.211.49
COPA 96.402.30 95.802.05 96.402.88 96.202.05 96.402.07 98.601.14 97.201.10 97.400.89 99.000.71 97.001.00
RTE 80.941.29 79.780.92 76.821.72 80.431.07 83.471.10 75.092.11 80.002.48 79.211.95 86.642.62 85.422.48
WiC 50.690.59 50.560.50 49.970.13 51.383.56 61.102.07 51.351.90 55.586.38 54.676.02 57.8711.08 64.838.59
ANLI-R1 46.244.21 42.543.20 40.263.66 40.286.46 59.820.51 63.062.63 71.923.71 73.563.88 72.308.05 75.003.03
ANLI-R2 40.440.90 38.360.82 38.443.46 41.884.50 50.160.82 58.401.19 65.363.75 65.481.91 64.982.94 67.302.34
ANLI-R3 42.530.99 38.781.04 43.675.25 37.500.81 55.751.66 61.353.14 67.320.98 66.230.72 63.036.03 66.380.74
Avg. 66.20 62.39 64.98 67.65 74.41 75.16 77.83 76.89 76.13 81.09
@@ 5.2 Main Experiments
{tblr}
width = column2-15 = c, vline2,15 = -, hline1-2,4 = -, hline1,4 = -0.08em, stretch = 0.5 Method & SST-2 MNLI QNLI MRPC QQP BoolQ CB COPA RTE WiC ANLIR1 ANLIR2 ANLIR3 Avg.
BC 95.4 75.0 83.5 68.6 70.3 87.9 75.0 98.0 84.1 63.3 59.8 51.1 53.3 74.3
BCL 96.3 75.0 83.5 74.3 72.3 88.8 83.9 99.082.763.258.049.752.2 75.3
@@
We further analyze the performance of BC by varying the ICL shots from 0 to 4 shots as shown in Fig. 6, and BC again outperforms all baseline methods. We also observe an overall trend for improved performance when more shots are available, and the performance disparities between BC and ICL converge on some tasks, which suggests that BC allows LLMs to more effectively take advantage of more in-context demonstrations. We also observe that PC exhibits the worst stability across the number of shots. In Table 3, we show that further slight gains, 1% on average, can be achieved by involving an adjustable strength parameter that refines the calibration and mitigates estimation errors. This alternative design not only makes BC applicable to both zero-shot and few-shot setups but also shows its capability to improve further from limited labeled data.
@@ 5.3 Robustness and Ablation Studies
We analyze the robustness of BC with respect to common prompt engineering design choices that were previously shown to significantly affect LLM performance (Lu et al., 2022; Liu et al., 2022a): choices and orders of in-context examples, the prompt template for ICL, and the verbalizers, as shown in Fig. 9 evaluated on RTE. Setup details are listed in Appendix §F. First, we find that BC is more robust to ICL choices and can mostly achieve the same performance with different ICL examples. Additionally, given a single set of ICL shots, altering the order between each ICL example has minimal impact on the BC performance. However, it is worth noting that an optimal order selection can still lead to promising ICL performance. Furthermore, we analyze the robustness of BC under 10 designs of prompt templates, where BC shows consistent improvement over the ICL baseline. Therefore, though BC makes further improvements, a well-designed template can further enhance the performance of BC. Lastly, we examine the robustness of BC to variations in verbalizer designs. Remarkably, even when employing unconventional choices such as emoji pairs as the verbalizers (listed in Tables 7 & 10) leading to dramatic oscillations of ICL performance, BC largely recovers performance. This observation shows BC robustifies LLM predictions under common prompt design choices and makes prompt engineering easier.
@@ Appendix A Additional Related Work
Prompt learning is an efficient learning pipeline for LLM as an alternative to traditional full-model fine-tuning (Liu et al., 2023). Soft prompting (Li & Liang, 2021; Lester et al., 2021; Liu et al., 2022b) enables fast adaptation of LLM by appending learnable continuous prompts in the embedding space while freezing the rest of weights. The recent development of parameter-efficient fine-tuning methods (Houlsby et al., 2019; Hu et al., 2022; Zhou et al., 2024), which learn additional modules, may also be interpreted as a form of soft prompting (He et al., 2022). However, these soft prompt learning methods inevitably require gradients and internal model access. On the other hand, hard prompting (Shin et al., 2020) is an appealing learning category for learning discrete prompts. Recent efforts have been devoted to black-box prompt search without accessing model gradients, and more interpretable prompts can be found by reinforcement learning (Deng et al., 2022; Zhang et al., 2023), gradient estimation (Diao et al., 2023), and other derivative-free search algorithms (Prasad et al., 2023; Zhou et al., 2023).
@@
Test-time adaptation aims to mitigate the covariate shift using the test-time statistics. Wang et al. propose to minimize the entropy by updating the affine parameters in the BN layer. Nado et al. (2020) and Schneider et al. (2020) introduce using test-time batch statistics for the standardization in BN and mixing it with source statistics to conquer covariate shift, respectively. Similarly, mixing the statistics with predefined hyperparameters (You et al., 2021; Khurana et al., 2021), interpolating source and target-domain statistics (Lim et al., 2023), or using a running average estimate (Mirza et al., 2022; Hu et al., 2021) have also been proposed to adapt the BN layer. Zou et al. (2022) introduce strength parameters in adapting the standardization statistics in semantic segmentation tasks. We differentiate from test-time BN approaches by mitigating the bias in the novel context of LLM, and there is no source statistic similar to a BN layer in computer vision backbones.
@@ Appendix C Dataset Statistics
{tblr}
cell23 = c, cell24 = c, cell25 = c, cell33 = c, cell34 = c, cell35 = c, cell43 = c, cell44 = c, cell45 = c, cell53 = c, cell54 = c, cell55 = c, cell63 = c, cell64 = c, cell65 = c, cell73 = c, cell74 = c, cell75 = c, cell83 = c, cell84 = c, cell85 = c, cell93 = c, cell94 = c, cell95 = c, cell103 = c, cell104 = c, cell105 = c, cell113 = c, cell114 = c, cell115 = c, cell123 = c, cell124 = c, cell125 = c, cell133 = c, cell134 = c, cell135 = c, cell143 = c, cell144 = c, cell145 = c, hline1-2,15 = -, hline1,15 = -0.08em, stretch=0.5 Dataset & Objective #sentences #classes |Test|